Write an equation for the line:
Locate the point where the line x + y= -5 and the line x- 2y =4 intersect. Write the equation of the line that goes through this point of intersection and the point (1/2, -2/5)

Answer by Alan3354(69443)

Locate the point where the line x + y= -5 and the line x- 2y =4 intersect.
x + y = -5
x -2y = 4
------------------ Subtract
3y = -9
y = -3
-------
x = -2
--> (-2,-3)
===================================
Write the equation of the line that goes through this point of intersection and the point (1/2,-2/5)
Find the slope m.
m = diffy/diffx = (13/5)/(5/2) = 26/25
Use y-y1 = m*(x-x1) and the point (-2,-3)
y + 3 = (26/25)*(x + 2)